OpenStudy (anonymous):

The U.S. Maritime Administration estimated that the cost per ton of building an oil tanker could be represented by the model f(x)=102,000/x+235 where y is the cost in dollars per ton and x is the tons (in thousands). What size of oil tanker (in thousands of tons) can be built for $350 per ton? a. 56 thousand tons b. 6 thousand tons c. 174 thousand tons d. 526 thousand tons

OpenStudy (saifoo.khan):

\[\frac{102000}{350+235}\]

OpenStudy (saifoo.khan):

c. 174 thousand tons

OpenStudy (anonymous):

saifoo.khan is my hero :)

OpenStudy (saifoo.khan):

hehe, thank you. :)

OpenStudy (anonymous):

do you know these?....If y varies inversely as x, find the inverse variation for the situation... y=65 when x=9 AND WHEN y=5 x=4

OpenStudy (saifoo.khan):

Yes, i know all these.

OpenStudy (anonymous):

y=65 x=9 not sure if it's y=585/x or 1/585x?... x/585? i don't know how to write it...

OpenStudy (anonymous):

i can figure them out, it's not knowing the format it should go in?

OpenStudy (saifoo.khan):

\[y = \frac k x\] \[65 = \frac k9 \] \[k = 65 \times 9 \] \[k = 585\]

OpenStudy (saifoo.khan):

So, \[y = \frac{585}{x}\]

OpenStudy (anonymous):

okay so the y=5 x= 4 would be y= 20/x ?

OpenStudy (saifoo.khan):

No, it actually depends on the formula they have given/

OpenStudy (anonymous):

Crap.. ok..

OpenStudy (saifoo.khan):

lol.

OpenStudy (saifoo.khan):

what have they given?

OpenStudy (anonymous):

It says if y varies inversely as x, find the inverse variation equation for the situation. y=5 x=4 a. y=5/4x b. y=20/x c. y=1/20x d. y= x/20

OpenStudy (saifoo.khan):

y = k/x 20 = k y = k/x

OpenStudy (saifoo.khan):

So, B.

OpenStudy (anonymous):

thanks again, for real. :)

OpenStudy (saifoo.khan):

You welcome..

OpenStudy (saifoo.khan):

Can i ask are you doing online college or something?

OpenStudy (anonymous):

Yeah ... and teaching myself math isn't that awesome.

OpenStudy (saifoo.khan):

Online college is easy. i know the math part. LOL.

OpenStudy (anonymous):

any other class i have noooo problem. but math, it's BAD lol :) you're very lucky it's easy for you!

OpenStudy (saifoo.khan):

lol, that's the case with mostly all of the Online school students.

OpenStudy (anonymous):

for me, it's literally like teaching myself again.. aaand i'm not a good teacher. and i have a 1 and 2 year old that don't help ;)

OpenStudy (saifoo.khan):

LOL. well if you want any help, you can contact me as well if you want.. =]

OpenStudy (anonymous):

thanks so much!!.. i mean that so sincerely! have a good day :D

OpenStudy (saifoo.khan):

uw. same 2 u. :]